Evaluate the determinate in the matrix.
|1 -3|
|-8 2|

Answer by Alan3354(69443)

Evaluate the determinate in the matrix.
|1 -3|
|-8 2|
----------
Det = 1*2 - (-8*-3) = 2 - 24
= -22